# Self-Supervised Multi-Part Articulated Objects Modeling via Deformable Gaussian Splatting and Progressive Primitive Segmentation

### v1

@@ Abstract
Articulated objects are ubiquitous in everyday life, and accurate 3D representations of their geometry and motion are critical for numerous applications. However, in the absence of human annotation, existing approaches still struggle to build a unified representation for objects that contain multiple movable parts. We introduce DeGSS, a unified framework that encodes articulated objects as deformable 3D Gaussian fields, embedding geometry, appearance, and motion in one compact representation. Each interaction state is modeled as a smooth deformation of a shared field, and the resulting deformation trajectories guide a progressive coarse-to-fine part segmentation that identifies distinct rigid components, all in an unsupervised manner. The refined field provides a spatially continuous, fully decoupled description of every part, supporting part-level reconstruction and precise modeling of their kinematic relationships. To evaluate generalization and realism, we enlarge the synthetic PartNet-Mobility benchmark and release RS-Art, a real-to-sim dataset that pairs RGB captures with accurately reverse-engineered 3D models. Extensive experiments demonstrate that our method outperforms existing methods in both accuracy and stability.
@@ 1 Introduction
Articulated objects, which range from hinged doors and sliding drawers to foldable laptops and many other everyday items with movable parts, are ubiquitous in both physical and virtual environments. Accurate modeling of their geometry and kinematics is indispensable for a wide spectrum of downstream tasks because it empowers robots to manipulate objects reliably [1, 2, 3, 4], enables plausible interactions in AR/VR scenes [5, 6, 7], and provides structured assets for digital twins [8, 9, 10]. These requirements motivate research on unified frameworks that simultaneously recover the geometry, appearance, and motion dynamics of articulated objects at the part level.
Earlier work on articulated object reconstruction can be divided into two strands. The first strand relies on explicit geometric representations and assumes access to clean CAD models with dense part annotations [11, 12], a requirement that limits applicability to real-world sensory data. The second strand adopts continuous implicit shape representations [13] such as signed distance fields to capture the deformation behavior [14, 15, 16, 17]. Their simplified appearance models and coarse geometric priors confine these methods to structurally simple objects with limited diversity.
Recent progress in radiance field rendering, including Neural Radiance Fields (NeRFs) [18] and 3D Gaussian Splatting (3DGS) [19], offers dense photometric self-supervision that jointly captures view-dependent appearance and geometry, providing a prior for articulated object modeling. PARIS [20] first employs a NeRF backbone [18] to analyze two multiview image sets captured in different interaction states and reconstructs separate static and dynamic fields for part-level modeling. Subsequent studies [21, 22] introduce 3DGS [19] to achieve higher rendering fidelity and faster convergence. However, these pipelines assume a single movable part observed in only two interaction states. As a result, each object is represented by one static and one moving component, which confines the framework to simple single-joint objects. Some approaches [23, 24] reconstruct radiance fields for two interaction states, compare the resulting geometry or fit a transform between the states, and then infer a multi-part decomposition. However, they require the number of parts to be known beforehand, and the reliance on exactly two discrete states forfeits any ability to represent continuous motion. In general, the above methods for modeling objects with multiple parts either tackle one joint at a time or assume a predefined part count, and both strategies require explicit correspondences between parts observed in different states. These post hoc correspondence procedures presume strong cross state consistency and prior structural knowledge, conditions that are difficult to satisfy in real scenarios.
Our key insight is that every interaction state of an articulated object can be described as a continuous deformation of a single standard 3D Gaussian field. Specifically, a latent-conditioned deformation network displaces each Gaussian primitive across states, unifying geometry, appearance, and kinematics in one representation and requiring only a handful of multiview image sets for supervision. Differences in deformation trajectories reveal the underlying rigid parts, which we segment through a progressive process: primitives are first clustered by trajectory similarity and then refined with boundary-aware Gaussian splitting that sharpens part interfaces. Drawing on this idea, we introduce DeGSS, a framework that couples Deformable 3D Gaussian Splatting with the Progressive coarse-to-fine Gaussian Primitives Segmentation strategy. Compared to recent methods [20, 23, 21, 22, 24], our approach handles articulated objects with multiple movable parts observed in more than two states, decouples the motion of each part, and smoothly extrapolates to unseen configurations, all without manual annotation.
Moreover, most previous studies [20, 23, 24, 22] test only a single object per category in PartNet-Mobility [1], a synthetic corpus that is too limited for rigorous evaluation and provides little insight into real-world performance. We enlarge the sample size per category in PartNet-Mobility and additionally release RS-Art, a benchmark that pairs multimodal recordings captured for real objects with their corresponding synthetic models, thereby narrowing the gap between simulation and reality.
@@
  • We propose DeGSS, a unified framework that represents articulated objects as deformable Gaussian fields and simultaneously recovers geometry, appearance, and kinematics without additional human priors.
  • We introduce a coarse-to-fine Gaussian segmentation procedure that groups Gaussians by trajectory similarity and refines the clusters, yielding accurate part-aware segmentation.
  • We extend existing benchmarks by enriching PartNet-Mobility and introducing RS-Art, a real-to-sim dataset, and validate the effectiveness and robustness of DeGSS across both benchmarks.
@@ 2 Related Work
Articulated object modeling.
Early PointNet pipelines predict part labels and motion from full scans, but depend on CAD-grade geometry and dense supervision, limiting their use in real settings [25, 26, 12]. Implicit SDF methods relax the input requirement, yet still need watertight, pre-aligned frames, a fixed joint budget, and usually only two interaction states [14, 16, 15]. Label-free NeRF approaches such as PARIS fit one static and one dynamic field to a single-joint object, and later 3D Gaussian-splatting variants improve visual fidelity while retaining the same “two-state, one-part” assumption [20, 21, 22]. Recent two-field registration schemes address multi-part objects by reconstructing each state separately and then aligning the fields, but they still require the part count beforehand and assume that the two states are closely aligned in configuration [23, 24]. DeGSS removes these restrictions: a latent-conditioned deformation field models any number of interaction states, while trajectory cues infer the part count automatically, yielding the first fully self-supervised framework for complex multi-part articulation.
Dynamic Gaussian splatting.
Recent work extends static 3D Gaussian splats by making the Gaussian parameters continuous functions of time. Dynamic 3D Gaussians [27] allow each primitive to translate and rotate through a time-conditioned MLP that enforces local rigidity. Four-dimensional Gaussian Splatting [28] embeds Gaussians in a joint space-time volume and renders any time stamp in real time. DynGS [29] introduces a learned trajectory basis that drives per-point motion with a scalar time code, and Explicit-4D Gaussian [30] adopts a similar design while regularizing appearance across frames. Later variants compress or distill these time-dependent fields to improve speed or attach semantics [31]. Although effective for smoothly varying videos, these methods encode dynamics with a single continuous clock, which makes it difficult to model discrete interaction states or multimodal motions where parts jump between configurations. Our method instead learns a conditioned deformation field in which every interaction state is indexed by an individual latent code, enabling explicit state-to-state mapping across arbitrarily many discrete states.
SAM-based 3D segmentation.
Recent efforts lift SAM [32] to neural field segmentation by projecting 2D masks or distilling them into Gaussian features. SA3D [33] propagates a single SAM mask through a trained NeRF to label voxels in three dimensions, and SANeRF HQ [34] improves boundary accuracy by aggregating view masks with density and color cues. Within the Gaussian domain, SAGA [35] learns per Gaussian embeddings under thousands of SAM masks to enable promptable instance grouping, whereas SAGD [36] decomposes boundary Gaussians before fusing multi-view masks for sharper labels. Gaussian Grouping [37] reconstructs scenes with 3DGS and directly lifts SAM masks to obtain object-level segmentation, and SAM3D [38] applies similar ideas to point clouds and meshes. All these methods view segmentation as a static task and rely on externally supplied two-dimensional prompts or per frame masks, so they cannot exploit how primitives move between interaction states. Our framework instead analyses Gaussian deformation trajectories to generate state-aware prompts automatically, then refines them with SAM guided splitting, which yields unmanual-prompt part segmentation.
@@
Figure 1: Overview of the DeGSS pipeline. Solid arrows indicate differentiable modules that participate in the joint optimization, whereas dashed arrows correspond to non-differentiable post-processing stages executed outside the optimization loop.
@@ 3 Methodology
This section explains the pipeline shown in Fig. 1. We first review the basics of 3D Gaussian Splatting (Sec. 3.1) and then attach a latent-conditioned deformation network that maps a standard Gaussian field to every interaction state (Sec. 3.2). On the resulting state-specific fields, we detect rigid parts from motion cues (Sec. 3.3) and refine their boundaries with visibility-guided SAM prompting followed by boundary-aware Gaussian splitting (Sec. 3.4). The refined part-aware Gaussian field then supports the multi-task modeling of each articulated object (Sec. 3.5).
@@ 3.1 Preliminaries: 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) [19] represents a scene by a finite set of oriented anisotropic Gaussian primitives
$\mathcal{G}=\bigl{\{}g_{i}=(\boldsymbol{\mu}_{i},\;\boldsymbol{\Sigma}_{i},\;\mathbf{c}_{i},\;\alpha_{i})\bigr{\}}_{i=1}^{N},$ (1)
where $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$ denotes the center of the $i$-th Gaussian, $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$ its full covariance that encodes both scale and orientation, $\mathbf{c}_{i}\in[0,1]^{3}$ the RGB color, and $\alpha_{i}\in[0,1]$ the opacity.
A perspective camera is modeled by a projection $\Pi:\mathbb{R}^{3}\!\to\!\mathbb{R}^{2}$. Linearization $\Pi$ at $\boldsymbol{\mu}_{i}$ maps the 3D covariance to a 2D screen-space covariance $\mathbf{S}_{i}\in\mathbb{R}^{2\times 2}$ and the mean pixel coordinates $\mathbf{u}_{i}=\Pi(\boldsymbol{\mu}_{i})$. Each primitive therefore induces the density
@@
All parameters $\{\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\mathbf{c}_{i},\alpha_{i}\}_{i=1}^{N}$ are optimized by minimizing a photometric loss between the rendered color $\mathbf{C}(\mathbf{x})$ and the corresponding ground-truth image.
@@ 3.2 Deformable Gaussian Splatting
To capture the full continuum of interaction states we train an implicit network, conditioned on a latent code, that warps a standard Gaussian field into state-specific configurations.
@@
Each Gaussian primitive is $g_{i}=(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\mathbf{c}_{i},\alpha_{i})$, with the covariance factorised as
$\boldsymbol{\Sigma}_{i}\;=\;\mathbf{R}\!\left(\mathbf{q}_{i}\right)\,\operatorname{Diag}\!\left(\mathbf{s}_{i}\right)\,\mathbf{R}\!\left(\mathbf{q}_{i}\right)^{\!\top},$ (4)
@@
where $\mathbf{R}(\boldsymbol{q}_{i})$ converts the unit quaternion $\boldsymbol{q}_{i}\in SO(3)$ into its $3\times 3$ rotation matrix, and $\mathrm{Diag}(\boldsymbol{s}_{i})$ places the positive scale vector $\boldsymbol{s}_{i}\in\mathbb{R}_{>0}^{3}$ on the diagonal of a $3\times 3$ matrix. This factorization separates orientation from per-axis stretch, keeping subsequent rotation and scale updates both intuitive and numerically stable.
@@
Given a latent code $\boldsymbol{\alpha}\!\in\!\mathbb{R}^{D}$ that encodes one interaction state, a multi-layer perceptron $f_{\text{def}}$ predicts per-primitive offsets
$(\Delta\boldsymbol{\mu}_{i},\Delta\mathbf{q}_{i},\Delta\mathbf{s}_{i})\;=\;f_{\text{def}}(\boldsymbol{\mu}_{i},\mathbf{q}_{i},\mathbf{s}_{i}\mid\boldsymbol{\alpha}),$ (5)
@@
with $\Delta\boldsymbol{\mu}_{i}\!\in\!\mathbb{R}^{3}$, $\Delta\boldsymbol{s}_{i}\!\in\!\mathbb{R}^{3}$, and an unconstrained four-vector $\Delta\boldsymbol{q}_{i}$ that is normalized to $SO(3)$ by $\Delta\boldsymbol{q}_{i}\leftarrow\Delta\boldsymbol{q}_{i}/\lVert\Delta\boldsymbol{q}_{i}\rVert_{2}$.
@@
Applying the learned offsets converts the standard Gaussian field into the configuration of a particular interaction state $k$:
$\boldsymbol{\mu}_{i}^{(k)}=\boldsymbol{\mu}_{i}+\Delta\boldsymbol{\mu}_{i},\qquad\mathbf{s}_{i}^{(k)}=\mathbf{s}_{i}+\Delta\mathbf{s}_{i},\qquad\mathbf{q}_{i}^{(k)}=\Delta\mathbf{q}_{i}\otimes\mathbf{q}_{i},$ (6)
@@
where $\otimes$ denotes multiplication of quaternions. Normalizing $\Delta\boldsymbol{q}_{i}$ produces a unit quaternion, so $\boldsymbol{q}_{i}^{(k)}$ remains a valid orientation. Eq. (6) therefore converts the canonical Gaussians $\{\boldsymbol{\mu}_{i},\boldsymbol{s}_{i},\boldsymbol{q}_{i}\}$ into the state-specific set $\{\boldsymbol{\mu}_{i}^{(k)},\boldsymbol{s}_{i}^{(k)},\boldsymbol{q}_{i}^{(k)}\}$ that describes the interaction state $k$.
@@
For the $K$ observed interaction states of the object, we jointly optimize the network parameters $\theta_{f_{\text{def}}}$, the latents $\{\boldsymbol{z}_{k}\}_{k=1}^{K}$, and the standard Gaussian parameters by minimizing the results.
@@
$\mathcal{L}_{\text{total}}=\underbrace{\sum_{k,v}\!\sum_{\mathbf{x}\in\Omega_{k}^{(v)}}\lVert\mathbf{C}_{k}^{(v)}(\mathbf{x})-\mathbf{I}_{k}^{(v)}(\mathbf{x})\rVert_{1}}_{\mathcal{L}_{\text{photo}}}+\mathcal{L}_{\text{D\_SIMM}},$ (7)
@@
where $\mathbf{C}_{k}^{(v)}$ is the color rendered in pixel $\mathbf{x}$, $\mathbf{I}_{k}^{(v)}$ is the corresponding ground truth image, and $\mathcal{L}_{\text{D\_SIMM}}$ is the standard 3DGS density similarity term that discourages overlapping Gaussians with dissimilar colors.
@@
After convergence, every latent $\boldsymbol{\alpha}_{k}$ produces a state-specific geometry $\{\boldsymbol{\mu}_{i}^{(k)},\boldsymbol{s}_{i}^{(k)},\boldsymbol{q}_{i}^{(k)}\}_{i=1}^{N}$ while sharing the appearance attributes $\{\mathbf{c}_{i},\alpha_{i}\}_{i=1}^{N}$. A single model therefore yields a coherent Gaussian scene for each interaction state and jointly encodes geometry, appearance, and articulation.

### v2

# PD2GS: Part-Level Decoupling and Continuous Deformation of Articulated Objects via Gaussian Splatting

@@ Abstract
Articulated objects are ubiquitous and important in robotics, AR/VR, and digital twins. Most self-supervised methods for articulated object modeling reconstruct discrete interaction states and relate them via cross-state geometric consistency, yielding representational fragmentation and drift that hinder smooth control of articulated configurations. We introduce PD2GS, a novel framework that learns a shared canonical Gaussian field and models the arbitrary interaction state as its continuous deformation, jointly encoding geometry and kinematics. By associating each interaction state with a latent code and refining part boundaries using generic vision priors, PD2GS enables accurate and reliable part-level decoupling while enforcing mutual exclusivity between parts and preserving scene-level coherence. This unified formulation supports part-aware reconstruction, fine-grained continuous control, and accurate kinematic modeling, all without manual supervision. To assess realism and generalization, we release RS-Art, a real-to-sim RGB-D dataset aligned with reverse-engineered 3D models, supporting real-world evaluation. Extensive experiments demonstrate that PD2GS surpasses prior methods in geometric and kinematic accuracy, and in consistency under continuous control, both on synthetic and real data.
Figure 1: Given multi-view image sets at arbitrary interaction states, our framework (a) builds a unified 3D Gaussian representation for articulated objects, (b) achieves part-level articulated object reconstruction and joint motion analysis, and (c) enables continuous deformation of articulated objects with multi-part decoupling by interpolating latent codes.
@@ 1 Introduction
Articulated objects, from hinged doors and sliding drawers to foldable laptops, are ubiquitous in physical and virtual environments. Accurate 3D representations of articulated objects are critical for robotics (Simeonov et al., 2022; Si et al., 2024; Zhao et al., 2024), AR/VR (Yang et al., 2024a; Mangalam et al., 2024), and digital twins (Li et al., 2025; Guo et al., 2024). Prior work on articulated object modeling often builds on access to annotated 3D models with dense part and kinematic labels (Wang et al., 2019; Yan et al., 2019; Mu et al., 2021; Jiang et al., 2022; Wei et al., 2022; Wang et al., 2024). Such strong supervision, together with simplified appearance models and coarse geometric priors, limits applicability to objects with simple kinematic structure and low part diversity.
Recent advances in neural rendering, including NeRF (Mildenhall et al., 2021) and 3DGS (Kerbl et al., 2023), enable multi-view learning of appearance and geometry, offering a foundation for annotation-efficient articulated object modeling. PARIS (Liu et al., 2023) first employs a NeRF backbone (Mildenhall et al., 2021) for part-level modeling across two distinct interaction states. Subsequent studies (Guo et al., 2025; Wu et al., 2025), incorporating 3DGS (Kerbl et al., 2023), improve rendering fidelity and convergence speed. However, these methods typically assume a single movable part across two states, limiting them to simple single-joint objects. In response, several approaches infer multi-part decompositions by contrasting two state-specific 3D representations (Weng et al., 2024; Liu et al., 2025b), typically via explicit Marching-Cubes meshes and per-state Gaussian fields with subsequent joint alignment. This discretization confines optimization to pairwise geometric matching between the two states, typically augmented with physics-based alignment constraints, which relies on exact correspondences and precludes modeling continuous part motion. Additionally, operating on exactly two interaction states provides only sparse observations of the articulation, which limits estimation accuracy and hampers the resolution of motion ambiguities. Recent dynamic variants of NeRF (Mildenhall et al., 2021) and 3DGS (Kerbl et al., 2023) typically learn continuous representations of dynamic scenes conditioned on time, enabling high-fidelity novel-view synthesis from temporally ordered inputs (Pumarola et al., 2021; Wu et al., 2022; Yang et al., 2024c; Luiten et al., 2024). However, articulated object modeling explicitly requires the decoupling of motions at the part level, whereas existing dynamic rendering methods inherently capture only holistic scene transformations. Moreover, without continuous temporal data, these dynamic approaches struggle to effectively disentangle individual part motions from a limited number of discrete interaction states, often resulting in severe geometric distortions and blurry rendering artifacts.
Our key insight is that each interaction state of an articulated object can be modeled as a continuous deformation of a shared canonical 3D Gaussian field, with per-primitive deformation trajectories coherent within parts and distinct across parts, thereby inducing part semantics and enabling unsupervised part-level grouping. Building on this idea, we introduce PD2GS, which realizes Part-Level Decoupling and Continuous Deformation via Gaussian Splatting, enabling part-aware reconstruction and smooth transitions to previously unseen configurations (Fig. 1). We design Deformable Gaussian Splatting that represents each interaction state with a latent code used to parameterize per-primitive transformations of a shared canonical Gaussian field. To enable part-level decoupling of Gaussian primitives, we propose a coarse-to-fine segmentation procedure. We first obtain a coarse grouping by clustering primitives according to deformation-trajectory similarity with guidance from a vision–language model. We then refine part boundaries via a boundary-aware splitting stage that employs a tailored 3D-2D prompting scheme for SAM, yielding sharp interfaces and fine-grained per-primitive labels while preserving smooth part motion. Accordingly, our framework operates entirely within the 3DGS optimization paradigm, is agnostic to the number of interaction states, and avoids imposing geometric or physics-based constraints that are not expressible within the Gaussian primitive parameterization.
Moreover, most prior studies (Liu et al., 2023; Weng et al., 2024; Liu et al., 2025b; Wu et al., 2025) evaluate at most one instance per category on PartNet-Mobility (Xiang et al., 2020), a synthetic corpus with limited intra-category diversity and weak evidence for real-world generalization. We enlarge per-category coverage on PartNet-Mobility and additionally release RS-Art, a real-to-sim evaluation dataset that pairs multi-view RGB captures of real objects with their reverse-engineered 3D models, enabling more rigorous assessment of sim-to-real performance.
@@
  • We introduce PD2GS, a self-supervised framework that learns a canonical Gaussian field and realizes interaction states as its continuous deformations, enabling part-level decoupling and the joint recovery of geometry, appearance, and kinematics.
  • We propose a coarse-to-fine segmentation of Gaussian primitives driven by deformation trajectories, with boundary refinement for sharp part interfaces and smooth part motion, yielding accurate part-aware segmentation.
  • We evaluate multiple instances per category and release RS-Art, a real-to-sim evaluation dataset pairing real RGB captures with reverse-engineered 3D models. Comprehensive experiments demonstrate that PD2GS outperforms prior methods.
@@ 2 Related Work
Articulated object modeling. Early methods like PointNet require dense supervision and CAD-quality inputs (Qi et al., 2017a; b; Yan et al., 2019). Implicit SDFs relax input needs but demand watertight meshes, pre-aligned frames, and usually two interaction states (Mu et al., 2021; Wei et al., 2022; Jiang et al., 2022). Label-free NeRF and 3DGS approaches (Liu et al., 2023; Guo et al., 2025; Wu et al., 2025) support only single-joint, two-state objects. Recent two-field registration methods handle multiple parts but still assume known part counts and similar state configs (Weng et al., 2024; Liu et al., 2025b). Distinct from these reconstruction tasks, conditional generative approaches (Lei et al., 2023; Liu et al., 2025a) leverage diffusion models to synthesize articulated shapes from sparse inputs. However, these methods depend on extensive labeled datasets and often hallucinate plausible geometries that deviate from the true instance structure. PD2GS overcomes these limitations with a latent-conditioned deformation field that supports arbitrary states and infers part structure from motion, enabling fully self-supervised multi-part articulation modeling.
Dynamic Gaussian splatting. Recent work extends static 3D Gaussians to dynamic scenes by parameterizing attributes as time-variant functions. Methods like Deformable 3DGS (Yang et al., 2024c; Luiten et al., 2024) employ time-conditioned MLPs to predict deformations, often enforcing physical priors. For efficient spacetime representation, approaches (Wu et al., 2024; Yang et al., 2024b) such as 4D-GS factorize the high-dimensional volume into compact feature planes to accelerate querying. Addressing the ambiguity of implicit deformations, recent works adopt explicit motion modeling via keyframes or neural flow fields for better tracking (Lee et al., 2024; Sun et al., 2025). In the context of articulated object modeling, these approaches inherently capture holistic scene dynamics, fundamentally lacking the ability to distinguish different motion semantics. Moreover, when applied to non-continuous interaction states, their reliance on temporal continuity leads to trajectory ambiguity and geometric artifacts. In contrast, PD2GS employs a latent-conditioned deformation field to explicitly model deformations across discrete interaction states, establishing a unified representation that facilitates subsequent fine-grained part decoupling and control.
SAM-based 3D segmentation. Recent efforts lift SAM (Kirillov et al., 2023) to neural field segmentation by projecting 2D masks or distilling them into Gaussian features. SA3D (Cen et al., 2023) and SANeRF HQ (Liu et al., 2024) propagate and aggregate SAM masks in NeRF using multi-view and density cues. Within the Gaussian domain, SAGA (Cen et al., 2025) and SAGD (Hu et al., 2024) learn instance-aware embeddings or refine boundaries, while Gaussian Grouping (Ye et al., 2024) and SAM3D (Yang et al., 2023) directly lift 2D masks to 3D reconstructions. These approaches treat segmentation as static and depend on external 2D prompts or per-frame masks, failing to leverage motion cues across states. Our framework instead analyzes Gaussian motion trajectories to auto-generate state-aware prompts, then refines segments via SAM-guided splitting, enabling fully automatic part segmentation.
@@
Figure 2: Overview of the PD2GS pipeline. Solid arrows indicate differentiable modules that participate in the joint optimization, whereas dashed arrows correspond to non-differentiable post-processing stages executed outside the optimization loop.
@@ 3 Methodology
This section explains the pipeline shown in Fig. 2. We first review the basics of 3D Gaussian Splatting (Sec. 3.1) and then attach a latent-conditioned deformation network that maps a canonical Gaussian field to every interaction state (Sec. 3.2). On the resulting state-specific fields, we detect rigid parts from motion cues (Sec. 3.3) and refine their boundaries with visibility-guided SAM prompting followed by boundary-aware Gaussian splitting (Sec. 3.4). The refined part-aware Gaussian field then supports the multi-task modeling of each articulated object (Sec. 3.5).
@@ 3.1 Preliminaries: 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) represents a scene by a finite set of oriented anisotropic Gaussian primitives
$\mathcal{G}=\bigl\{g_{i}=(\bm{\mu}_{i},\;\bm{\Sigma}_{i},\;\mathbf{c}_{i},\;\alpha_{i})\bigr\}_{i=1}^{N},$ (1)
where $\bm{\mu}_{i}\in\mathbb{R}^{3}$ denotes the center of the $i$-th Gaussian, $\bm{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$ its full covariance that encodes both scale and orientation, $\mathbf{c}_{i}\in[0,1]^{3}$ the RGB color, and $\alpha_{i}\in[0,1]$ the opacity.
A perspective camera is modeled by a projection $\Pi:\mathbb{R}^{3}\!\to\!\mathbb{R}^{2}$. Linearization $\Pi$ at $\bm{\mu}_{i}$ maps the 3D covariance to a 2D screen-space covariance $\mathbf{S}_{i}\in\mathbb{R}^{2\times 2}$ and the mean pixel coordinates $\mathbf{u}_{i}=\Pi(\bm{\mu}_{i})$. Each primitive therefore induces the density
@@
All parameters $\{\bm{\mu}_{i},\bm{\Sigma}_{i},\mathbf{c}_{i},\alpha_{i}\}_{i=1}^{N}$ are optimized by minimizing a photometric loss between the rendered color $\mathbf{C}(\mathbf{x})$ and the corresponding ground-truth image.
@@ 3.2 Deformable Gaussian Splatting
To capture the full continuum of interaction states we train an implicit network, conditioned on a latent code, that warps a canonical Gaussian field into state-specific configurations.
@@
Each Gaussian primitive is $g_{i}=(\bm{\mu}_{i},\bm{\Sigma}_{i},\mathbf{c}_{i},\alpha_{i})$, with the covariance factorised as
$\bm{\Sigma}_{i}\;=\;\mathbf{R}\!\left(\mathbf{q}_{i}\right)\,\operatorname{Diag}\!\left(\mathbf{s}_{i}\right)\,\mathbf{R}\!\left(\mathbf{q}_{i}\right)^{\!\top},$ (4)
@@
where $\mathbf{R}(\bm{q}_{i})$ converts the unit quaternion $\bm{q}_{i}\in SO(3)$ into its $3\times 3$ rotation matrix, and $\mathrm{Diag}(\bm{s}_{i})$ places the positive scale vector $\bm{s}_{i}\in\mathbb{R}_{>0}^{3}$ on the diagonal of a $3\times 3$ matrix. This factorization separates orientation from per-axis stretch, keeping subsequent rotation and scale updates both intuitive and numerically stable.
@@
Given a latent code $\bm{\alpha}\!\in\!\mathbb{R}^{D}$ that encodes one interaction state, a multi-layer perceptron $f_{\text{def}}$ predicts per-primitive offsets
$(\Delta\bm{\mu}_{i},\Delta\mathbf{q}_{i},\Delta\mathbf{s}_{i})\;=\;f_{\text{def}}(\bm{\mu}_{i},\mathbf{q}_{i},\mathbf{s}_{i}\mid\bm{\alpha}),$ (5)
@@
with $\Delta\bm{\mu}_{i}\!\in\!\mathbb{R}^{3}$, $\Delta\bm{s}_{i}\!\in\!\mathbb{R}^{3}$, and an unconstrained four-vector $\Delta\bm{q}_{i}$ that is normalized to $SO(3)$ by $\Delta\bm{q}_{i}\leftarrow\Delta\bm{q}_{i}/\lVert\Delta\bm{q}_{i}\rVert_{2}$.
@@
Applying the learned offsets converts the canonical Gaussian field into the configuration of a particular interaction state $k$:
$\bm{\mu}_{i}^{(k)}=\bm{\mu}_{i}+\Delta\bm{\mu}_{i},\qquad\mathbf{s}_{i}^{(k)}=\mathbf{s}_{i}+\Delta\mathbf{s}_{i},\qquad\mathbf{q}_{i}^{(k)}=\Delta\mathbf{q}_{i}\otimes\mathbf{q}_{i},$ (6)
@@
where $\otimes$ denotes multiplication of quaternions. Normalizing $\Delta\bm{q}_{i}$ produces a unit quaternion, so $\bm{q}_{i}^{(k)}$ remains a valid orientation. Eq. 6 therefore converts the canonical Gaussians $\{\bm{\mu}_{i},\bm{s}_{i},\bm{q}_{i}\}$ into the state-specific set $\{\bm{\mu}_{i}^{(k)},\bm{s}_{i}^{(k)},\bm{q}_{i}^{(k)}\}$ that describes the interaction state $k$.
@@
For the $K$ observed interaction states of the object, we jointly optimize the network parameters $\theta_{f_{\text{def}}}$, the latents $\{{\bm{\alpha}}_{k}\}_{k=1}^{K}$, and the canonical Gaussian parameters by minimizing the following total loss:
@@
$\mathcal{L}_{\text{total}}=\underbrace{\sum_{k,v}\!\sum_{\mathbf{x}\in\Omega_{k}^{(v)}}\lVert\mathbf{C}_{k}^{(v)}(\mathbf{x})-\mathbf{I}_{k}^{(v)}(\mathbf{x})\rVert_{1}}_{\mathcal{L}_{\text{photo}}}+\mathcal{L}_{\text{D}_{\text{{SIMM}}}},$ (7)
@@
where $\mathbf{C}_{k}^{(v)}$ is the color rendered in pixel $\mathbf{x}$, $\mathbf{I}_{k}^{(v)}$ is the corresponding ground truth image, and $\mathcal{L}_{\text{D}_{\text{{SIMM}}}}$ is the canonical 3DGS density similarity term that discourages overlapping Gaussians with dissimilar colors.
@@
After convergence, every latent $\bm{\alpha}_{k}$ produces a state-specific geometry $\{\bm{\mu}_{i}^{(k)},\bm{s}_{i}^{(k)},\bm{q}_{i}^{(k)}\}_{i=1}^{N}$ while sharing the appearance attributes $\{\mathbf{c}_{i},\alpha_{i}\}_{i=1}^{N}$. A single model therefore yields a coherent Gaussian scene for each interaction state and jointly encodes geometry, appearance, and articulation.